Question 1) A square flower bed in Michelle's garden is to be enlarged by adding 3 feet on each side. If the larger square has an area of 169 square feet, what is the length of the original square?
Answer by ankor@dixie-net.com(22740)

A square flower bed in Michelle's garden is to be enlarged by adding 3 feet on
each side. If the larger square has an area of 169 square feet, what is the
length of the original square?
:
Let s = length of the side of the original square
then
(s+6) = length of the enlarged square when 3' is added to each side
:
Area of new square given as 169 sq/ft
(s+6)^2 = 169
FOIL
s^2 + 12s + 36 = 169
:
s^2 + 12s + 36 - 169 = 0
:
s^2 + 12s - 133 = 0
Factor this to
(s-7)(s+19) = 0
Positive solution
s = 7' is the side of the original square
:
;
Check solution:
(7+6)^2 = new area
13^2 = 169